OpenStudy (anonymous):

Solve the absolute value equation, if possible. If there is no solution, explain why. | 5x + 13| = -7

OpenStudy (anonymous):

solve like normal, and then take the absolute value of whatever x is. So if it's a negative just make it a positive basically

OpenStudy (dtan5457):

Absolute value equations, 2 solutions 5x+13=-7 5x+13=7 solve for both

OpenStudy (anonymous):

5x=20

OpenStudy (anonymous):

5/20=4

OpenStudy (anonymous):

is that right

OpenStudy (dtan5457):

none of them have 20 lol

OpenStudy (anonymous):

then how to do it

OpenStudy (perl):

hint: an absolute value expression can never be negative

OpenStudy (anonymous):

so I should change -7 to 7 or what

OpenStudy (jhannybean):

|5x+13|=-7 -(5x+13) = -7 5x+13 = -7

OpenStudy (jhannybean):

Solve for both.

OpenStudy (perl):

you're done. there's no solution

OpenStudy (perl):

The solution set is empty (thats a fancy way of saying no solutions).

OpenStudy (tkhunny):

No, no, no, no, no!!! |Stuff| >= 0 It can NEVER be -7. Just glance at it and state that and move on. There is no solution. No need to tinker with it!

OpenStudy (anonymous):

oh ok thanks

OpenStudy (perl):

You can also graph the right side and left side as separate functions. One can see there is no intersection . Note that you have to zoom out to see both curves. https://www.desmos.com/calculator/qdf6wzh7rs

OpenStudy (perl):

oops, typo in function https://www.desmos.com/calculator/bqbsjiytma

OpenStudy (anonymous):

oh ok